# Extremal correlation coefficient for functional data

### v2

@@ Abstract
We propose a coefficient that measures extremal dependence in paired samples of functions. It has properties similar to the Pearson correlation, but differs in significant ways: 1) it is designed to measure dependence between curves, 2) it focuses only on extreme curves. The new coefficient is derived within the framework of regular variation in Banach spaces. A consistent estimator is proposed and justified by an asymptotic analysis and a simulation study. The usefulness of the new coefficient is illustrated using financial and climate functional data.
@@ 1 Introduction
With the growing impact of extreme events such as financial downturns or unusual weather, there has been increasing interest in developing statistical tools to study patterns of extreme curves. This is to a large extent due to the increasing availability of high resolution data; asset price curves can be constructed at any temporal resolution, and modern weather databases and computer models contain measurements at hourly or even higher frequencies. Such data can be interpreted as curves, e.g., one curve per day, providing more comprehensive view of daily patterns compared to a single summary number like the closing price or maximum temperature. Analyzing extreme curves in the framework of functional data analysis thus leads to a more precise understanding of the impacts associated with extreme events.
This paper makes a methodological and theoretical contribution at the nexus of extreme value theory and functional data analysis. We propose a coefficient that quantifies the tendency of paired extreme curves to exhibit similar patterns simultaneously. Two examples of the type of questions that the tool deals with are the following: 1) During a stock market crisis, such as the market decline due to the COVID-19 pandemic, do returns of different sectors of the economy exhibit similar extreme daily trajectories? 2) How likely is location A to experience a similar daily pattern of temperature as location B (on the same day) during a heat wave? Our proposed coefficient offers a more precise quantification of extreme risk by focusing on 1) the shape of curves and 2) the extreme parts of paired samples. This point is further illustrated with a data example in Section 6.1.
@@
We note that there has been considerable research aimed at quantifying extremal dependence for heavy-tailed random vectors. Ledford and Tawn (1996, 1997, 2003) introduced the coefficient of tail dependence, which was later generalized to the extremogram by Davis & Mikosch (2009). The extremal dependence measure based on the angular measure of a regularly varying random vector was introduced by Resnick (2004) and further investigated by Larsson & Resnick (2012). Janßen et al. (2023) recently introduced a unified approach for representing tail dependence using random exceedence sets. Those measures for extremes are designed for random vectors in a Euclidean space. Therefore, applying any such measures to functional data requires some sort of dimension reduction, e.g., principal component analysis, or data compression like converting daily temperature curves to daily average or maximum values. The reduced data are then analyzed using those tools for multivariate extremes, see, e.g., Meinguet (2010)Dombry & Ribatet (2015), and Kim & Kokoszka (2022). This approach is convenient, but it does not fully utilize all relevant information that functional data contain.
@@
We develop a new measure, the extremal correlation coefficient, that captures the extremal dependence of paired functional samples utilizing the information in the sizes and shapes of the curves. The measure involves an inner product of pairs of extreme curves, and therefore requires finite second moment. Similar ideas have been applied in non-extreme contexts of functional data analysis. Dubin & Müller (2005) introduced a measure, called dynamical correlation, that computes the inner product of all pairs of standardized curves. The concept was further studied by Yeh et al. (2023) where an autocorrelation measure, termed spherical autocorrelation, for functional time series was proposed. These measures are however computed based on the total body of functional data, and so are not suitable for describing extremal dependence.
@@
The coefficient we develop quantifies extremal dependence by specifically focusing on the extreme parts of heavy-tailed functional observations. It is conceptually appealing, as it shares desirable features with the classic correlation coefficient: 1) its values range from -1 to 1, 2) it measures the strength and direction of linear relationship between two extreme curves, 3) if the extremal behavior of two curves is independent, the coefficient is zero. Moreover, it can be used in practice with a relatively simple numerical implementation. We thus hope that such interpretable and tractable tool makes a useful contribution.
@@
Turning to mathematical challenges, the concept of vague convergence, see e.g., Chapters 2 and 6 of Resnick (2007), cannot be readily used. The vague convergence, which now provides a standard mathematical framework for extremes in Euclidean spaces, can be defined only on locally compact spaces. Since every locally compact Banach space has finite dimension, a different framework must be used for functional data in Hilbert spaces. We use the theory of regularly varying measures developed by Hult & Lindskog (2006) who introduced the notion of $M_{0}$ convergence, which works for regularly varying measures on complete separable metric spaces. The $M_{0}$ convergence is further studied by Meinguet (2010), where it is applied to regularly varying time series in a separable Banach space. The concept of $M_{0}$ convergence has been generalized, see Section B.1.1. of Kulik & Soulier (2020), with notable contributions from Lindskog et al. (2014) and Segers et al. (2017). We establish the consistency of the estimator we propose within the framework developed by Hult & Lindskog (2006) and  Meinguet (2010). We proceed through a number of $M_{0}$ convergence results that allow us to apply an abstract Bernstein-type inequality. A method of computing the extremal correlation coefficients analytically (in relatively simple cases) is also developed.
The remainder of the paper is organized as follows. In Section 2, we review regularly varying random elements in Banach spaces. In Section 3, we extend the concept of regular variation to bivariate random elements in a product Banach space. The extremal correlation coefficient is introduced in Section 4, where its asymptotic properties are also studied. Section 5 contains a simulation study, and Section 6 illustrates applications to intraday return curves and daily temperature curves.
Theoretical justification of our approach requires more detailed background and some technical derivations, which are placed in the Supplementary material. Preliminary results for the proof of Theorem 4.4 are presented in Section A, followed by its proof in Section B. In Section C, we provide the proof of Lemma 5.1, and additional numerical results from Sections 5 and 6 are included in Sections D and E, respectively.
@@ 2 Regular variation in Banach spaces
We now briefly review the $M_{0}$ convergence in a separable Banach space ${\mathbb{B}}$. In what follows, ${0}$ is the zero element. Fix a norm $\|\cdot\|_{{\mathbb{B}}}$ and let $B_{\varepsilon}:=\{z\in{\mathbb{B}}:\|z\|_{{\mathbb{B}}}<\varepsilon\}$ be the open ball of radius $\varepsilon>0$ centered at the origin. A Borel measure $\mu$ defined on ${\mathbb{B}}_{0}:={\mathbb{B}}\setminus\{\mbox{\boldmath${0}$}\}$ is said to be boundedly finite if $\mu(A)<\infty$, for all Borel sets that are bounded away from ${0}$, i.e., $A\cap B_{\varepsilon}=\emptyset$, for some $\varepsilon>0$. Let $M_{0}({\mathbb{B}})$ be the collection of all such measures on ${\mathbb{B}}_{0}$. For $\mu_{n},\mu\in M_{0}({\mathbb{B}})$, the sequence of $\mu_{n}$ converges to $\mu$ in the $M_{0}$ topology ($\mu_{n}\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu$), if $\mu_{n}(A)\to\mu(A)$, for all bounded away from ${0}$, $\mu$–continuity Borel sets $A$, i.e., those with $\mu(\partial A)=0$, where $\partial A$ is the boundary of $A$. Equivalently, $\mu_{n}\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu$, if $\int_{{\mathbb{B}}}f(x)\mu_{n}(dx)\to\int_{{\mathbb{B}}}f(x)\mu(dx)$ for all $f\in{\mathcal{C}}_{0}({\mathbb{B}})$, where ${\mathcal{C}}_{0}({\mathbb{B}})$ is the class of bounded and continuous functions $f:{\mathbb{B}}_{0}\to{\mathbb{R}}$ that vanish on a neighborhood of ${0}$.
@@
A random element $X$ in ${\mathbb{B}}$ is regularly varying with index $-\alpha$, $\alpha>0$, if there exist a sequence $b(n)\to\infty$ and a measure $\mu$ in $M_{0}({\mathbb{B}})$ such that
$n{\rm pr}\left(\frac{X}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu,\ \ \ \ n\to\infty,$ (1)
where the exponent measure $\mu$ satisfies $\mu(tA)=t^{-\alpha}\mu(A)$ for Borel sets $A\subset{\mathbb{B}}_{0}$.
A possible choice for $b(n)$ is the quantile function, defined by ${\rm pr}(\|X\|_{{\mathbb{B}}}>b(n))=n^{-1}$. Roughly speaking, the tail probability of $X$ decays like a power function, ${\rm pr}(\|X\|_{{\mathbb{B}}}>t)\approx Ct^{-\alpha}$, as $t\to\infty$. The following lemma, see Hult & Lindskog (2006), states an equivalent definition of a regularly varying element in ${\mathbb{B}}$.
@@
A random element $X$ in ${\mathbb{B}}$ is regularly varying with index $-\alpha$, $\alpha>0$, if and only if there exist a sequence $b^{\prime}(n)\to\infty$ and a probability measure $\Gamma$ on ${\mathbb{S}}:=\{x\in{\mathbb{B}}:\|x\|_{{\mathbb{B}}}=1\}$ (called the angular measure) such that for any $y>0$,
$n{\rm pr}\left(\|X\|_{{\mathbb{B}}}>b^{\prime}(n)y,X/\|X\|_{{\mathbb{B}}}\in\cdot\right)\stackrel{{\scriptstyle w}}{{\longrightarrow}}cy^{-\alpha}\Gamma,\ \ \ \ n\to\infty,$ (2)
@@
If Definition 2.1 (or condition (2)) holds, we write $X\in RV(-\alpha,\Gamma)$. The polar representation (2) provides an intuitive interpretation of regular variation in ${\mathbb{B}}$. It characterizes regular variation of $X$ in ${\mathbb{B}}$ using two components, the tail index $\alpha$ and the angular probability measure $\Gamma$. The tail index $\alpha$ quantifies how heavy the tail distribution of $\|X\|_{{\mathbb{B}}}$ is, e.g., the probability of extreme curves occurring gets higher as $\alpha$ gets smaller. While the tail index $\alpha$ determines the frequency of occurrence of extreme curves, the angular measure $\Gamma$, defined on the unit sphere ${\mathbb{S}}$, fully characterizes the distribution of the shape of the scaled extreme curves, $X/\|X\|_{{\mathbb{B}}}$. To illustrate this, consider a set of orthonormal functions in $L^{2}([0,1])$ of the form
$\phi_{j}(t)=\sqrt{2}\sin\left(\left(j-\frac{1}{2}\right)\pi t\right),\ \ \ \ j=1,2,\ldots,\ \ t\in[0,1].$ (3)
The first three functions are shown in the left-most plot of Figure 1. We consider a finite-dimensional subspace of $L^{2}([0,1])$, spanned by the first 9 $\phi_{j}$’s, for the purpose of simulations. The data generating process is $X(t)=\sum_{j=1}^{9}Z_{j}\phi_{j}(t)$, where $[Z_{1},\ldots,Z_{9}]^{\top}$ is a 9-dimensional random vector with independent components. Suppose that $Z$ is a random variable following a Pareto distribution with tail index $\alpha=3$ and $N$ is a normal random variable with mean 0 and variance 0.5. We consider the following three cases for $[Z_{1},\ldots,Z_{9}]^{\top}$:
1. $[Z,N,N,N,\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{1}$.
2. $[N,Z,N,N\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{2}$.
3. $[N,N,Z,N\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{3}$.
In all three cases, it follows from Proposition 7.1 and Example 7.3 of Meinguet & Segers (2010) that $X(t)$ is regularly varying with tail index $\alpha=3$. Figure 1 displays simulated data with sample size of 100 for each of the three cases. The plots of simulated data clearly show that the angular measure $\Gamma$ represents the distribution of the shapes of extreme curves in that they are dominated by the shape of the functional axis $\phi_{j}$ on which $\Gamma$ concentrates.
@@ 3 Bivariate regular variation in Banach spaces
In order to describe the extremal dependence of two regularly varying random elements $X$ and $Y$ in $L^{2}$, we need to identify their joint probabilistic behavior. We again study it in the more general space ${\mathbb{B}}^{2}$. We propose the following definition.
@@
A bivariate random element $[X,Y]^{\top}$ in ${\mathbb{B}}^{2}$ is said to be jointly regularly varying with index $-\alpha$, $\alpha>0$, if there exist a sequence $b(n)\to\infty$ and a measure $\mu$ in $M_{0}({\mathbb{B}}^{2})$ such that
$n{\rm pr}\left(\frac{(X,Y)}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu,\ \ \ \ n\to\infty,$ (4)
where the joint exponent measure $\mu$ satisfies $\mu(tA)=t^{-\alpha}\mu(A)$ for Borel sets $A\subset{\mathbb{B}}_{0}^{2}$.
We assume that one-dimensional marginal distributions of $\mu$ are non-degenerate, i.e., $\mu_{X}:=\mu(\cdot\times{\mathbb{B}})$ and $\mu_{Y}:=\mu({\mathbb{B}}\times\cdot)$ are measures in $M_{0}({\mathbb{B}})$ satisfying analogs of (1). Since $X$ and $Y$ are normalized by the same function $b(n)$, the marginal distributions are tail equivalent. A possible choice for $b(n)$ is the quantile function, defined by
${\rm pr}(\|(X,Y)\|_{{\mathbb{B}}^{2}}>b(n))=n^{-1}.$
With this choice, we have that
$n{\rm pr}\left(\frac{(X,Y)}{b(n)}\in{\mathcal{A}}_{1}\right)=\frac{{\rm pr}\left((X,Y)\in b(n){\mathcal{A}}_{1}\right)}{{\rm pr}(\|(X,Y)\|_{{\mathbb{B}}^{2}}>b(n))}=\frac{{\rm pr}\left((X,Y)\in{\mathcal{A}}_{b(n)}\right)}{{\rm pr}\left((X,Y)\in{\mathcal{A}}_{b(n)}\right)}=1,$
where ${\mathcal{A}}_{r}$ is defined by
@@
Thus, it follows from the $M_{0}$ convergence in (4) and Lemma A.1 that
@@
which implies that $\mu$ is a probability measure on ${\mathcal{A}}_{1}$. Throughout the paper, we set

### v3

@@ Abstract
We propose a coefficient that measures extremal dependence in paired samples of functions. It has properties similar to the Pearson correlation, but differs in significant ways: (i) it is designed to measure dependence between curves, (ii) it focuses only on extreme curves. The new coefficient is derived within the framework of regular variation in Banach spaces. A consistent estimator is proposed and justified by an asymptotic analysis and a simulation study. The usefulness of the new coefficient is illustrated using financial and climate functional data.
@@ 1 Introduction
With the growing impact of extreme events such as financial downturns or unusual weather, there has been increasing interest in developing statistical tools to study patterns of extreme curves. This is to a large extent due to the increasing availability of high resolution data; asset price curves can be constructed at any temporal resolution, and modern weather databases and computer models contain measurements at hourly or even higher frequencies. Such data can be interpreted as curves, e.g., one curve per day, providing a more comprehensive view of daily patterns compared to a single summary number like the closing price or maximum temperature. Analyzing extreme curves in the framework of functional data analysis thus leads to a more precise understanding of the impacts associated with extreme events.
This paper makes a methodological and theoretical contribution at the nexus of extreme value theory and functional data analysis. We propose a coefficient that quantifies the tendency of paired extreme curves to exhibit similar patterns simultaneously. Two examples of the type of questions that the tool deals with are the following: (i) during a stock market crisis, such as the market decline due to the covid-19 pandemic, do returns of different sectors of the economy exhibit similar extreme daily trajectories? (ii) how likely is location A to experience a similar daily pattern of temperature as location B (on the same day) during a heat wave? Our proposed coefficient offers a more precise quantification of extreme risk by focusing on (i) the shape of curves and (ii) the extreme parts of paired samples. This point is further illustrated with a data example in Section 6.1.
@@
In the context of heavy-tailed random vectors, there has been considerable research aimed at quantifying extremal dependence. Ledford and Tawn (1996, 1997, 2003) introduced the coefficient of tail dependence, which was later generalized to the extremogram by Davis & Mikosch (2009). The extremal dependence measure based on the angular measure of a regularly varying random vector was introduced by Resnick (2004) and further investigated by Larsson & Resnick (2012). Janßen et al. (2023) recently introduced a unified approach for representing tail dependence using random exceedance sets. Those measures for extremes are designed for random vectors in a Euclidean space. Therefore, applying any such measures to functional data requires some sort of dimension reduction, e.g., principal component analysis, or data compression like converting daily temperature curves to daily average or maximum values. The reduced data are then analyzed using those tools for multivariate extremes, see, e.g., Meinguet (2010)Dombry & Ribatet (2015), and Kim & Kokoszka (2022). This approach is convenient, but it does not fully utilize all relevant information that functional data contain.
@@
We develop a new measure, the extremal correlation coefficient, that captures the extremal dependence of paired functional samples utilizing the information in the sizes and shapes of the curves. The measure involves the inner product of pairs of extreme curves and therefore requires a finite second moment. Similar ideas have been applied in non-extreme contexts of functional data analysis. Dubin & Müller (2005) introduced a measure, called dynamical correlation, that computes the inner product of all pairs of standardized curves. The concept was further studied by Yeh et al. (2023) where an autocorrelation measure, termed spherical autocorrelation, for functional time series was proposed. These measures are however computed based on the total body of functional data and so are not suitable for describing extremal dependence.
@@
The coefficient we develop quantifies extremal dependence by specifically focusing on the extreme parts of heavy-tailed functional observations. It is conceptually appealing, as it shares desirable features with the classic correlation coefficient: (i) its values range from -1 to 1, (ii) it measures the strength and direction of linear relationship between two extreme curves, (iii) if the extremal behavior of two curves is independent, the coefficient is zero. Moreover, it can be used in practice with a relatively simple numerical implementation. We thus hope that such an interpretable and tractable tool makes a useful contribution.
@@
Turning to mathematical challenges, the concept of vague convergence, see e.g., Chapters 2 and 6 of Resnick (2007), cannot be readily used. The vague convergence, which now provides a standard mathematical framework for extremes in Euclidean spaces, can be defined only on locally compact spaces. Since every locally compact Banach space is finite-dimensional, a different framework must be used for functional data in Hilbert spaces. We use the theory of regularly varying measures developed by Hult & Lindskog (2006) who introduced the notion of $M_{0}$ convergence, which works for regularly varying measures on complete separable metric spaces. The $M_{0}$ convergence is further studied by Meinguet (2010), where it is applied to regularly varying time series in a separable Banach space. The concept of $M_{0}$ convergence has been generalized, see Section B.1.1. of Kulik & Soulier (2020), with notable contributions from Lindskog et al. (2014) and Segers et al. (2017). We establish the consistency of the estimator we propose within the framework developed by Hult & Lindskog (2006) and Meinguet (2010). We proceed through a number of $M_{0}$ convergence results that allow us to apply an abstract Bernstein-type inequality. A method for analytically computing the extremal correlation coefficients in relatively simple cases is also developed.
@@ 2 Regular variation in Banach spaces
This section presents background needed to understand the development in Sections 3 and 4. In functional data analysis, observations are typically treated as elements of $L^{2}:=L^{2}({\mathcal{T}})$, where the measure space $\mathcal{T}$} is such that $L^{2}({\mathcal{T}})$ is a separable Hilbert space, equipped with the usual inner product $\left\langle x,y\right\rangle=\int_{{\mathcal{T}}}x(t)y(t)dt$. The $L^{2}$-norm is then $\|x\|=\left\langle x,x\right\rangle^{1/2}$ = $(\int_{{\mathcal{T}}}x(t)^{2}dt)^{1/2}$. An introduction to functional data analysis is presented in Kokoszka & Reimherr (2017), with a detailed mathematical treatment available in Hsing & Eubank (2015). While we refer to the elements of $L^{2}$ as curves, due to the examples we consider, the set $\mathcal{T}$} can be a fairly abstract space (a metric Polish space), for example a spatial domain.
@@
A random element $X$ in $\mathbb{B}$} is regularly varying with index $\alpha$, $\alpha>0$, if there exist a sequence $b(n)\to\infty$ and a measure $\mu$ in $M_{0}({\mathbb{B}})$ such that
$n{\rm pr}\left(\frac{X}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu,\ \ \ \ n\to\infty.$ (1)
The exponent measure $\mu$ is defined up to a multiplicative constant and satisfies $\mu(tA)=t^{-\alpha}\mu(A)$ for any $t>0$ and Borel sets $A\subset{\mathbb{B}}_{0}$. The following lemma, see Chapter 2 of Meinguet (2010), states an equivalent definition of regularly varying elements in $\mathbb{B}$}. Throughout the paper, we denote the unit sphere in a normed space $\mathbb{E}$} by
${\mathbb{S}}_{{\mathbb{E}}}:=\{u\in{\mathbb{E}}:\|u\|_{{\mathbb{E}}}=1\}.$
@@
A random element $X$ in $\mathbb{B}$} is regularly varying with index $\alpha$, $\alpha>0$, if and only if there exist a sequence $b^{\prime}(n)\to\infty$ and a probability measure $\Gamma$ on ${\mathbb{S}}_{{\mathbb{B}}}$ (called the angular measure) such that for any $z>0$,
$n{\rm pr}\left(\|X\|_{{\mathbb{B}}}>b^{\prime}(n)z,X/\|X\|_{{\mathbb{B}}}\in\cdot\right)\stackrel{{\scriptstyle w}}{{\longrightarrow}}cz^{-\alpha}\Gamma,\ \ \ \ n\to\infty,$ (2)
@@
If Definition 2.1 (or condition (2)) holds, we write $X\in RV(-\alpha,\Gamma)$. The polar representation (2) provides an intuitive interpretation of regular variation in $\mathbb{B}$}. It characterizes regular variation of $X$ in $\mathbb{B}$} using two components, the tail index $\alpha$ and the angular probability measure $\Gamma$. The tail index $\alpha$ quantifies the tail heaviness of the distribution of $\|X\|_{{\mathbb{B}}}$, with smaller values of $\alpha$ implying a higher probability of extreme curves. While the tail index $\alpha$ determines the frequency of occurrence of extreme curves, the angular measure $\Gamma$, defined on the unit sphere ${\mathbb{S}}_{{\mathbb{B}}}$, fully characterizes the distribution of the shape of the scaled extreme curves, $X/\|X\|_{{\mathbb{B}}}$. To illustrate this, consider a set of orthonormal functions in $L^{2}([0,1])$ of the form
$\phi_{j}(t)=2^{1/2}\sin\left(\left(j-\frac{1}{2}\right)\pi t\right),\ \ \ \ j=1,2,\ldots,\ \ t\in[0,1].$ (3)
The first three functions are shown in the left-most plot of Figure 1. We consider a finite-dimensional subspace of $L^{2}([0,1])$, spanned by the first 9 $\phi_{j}$’s, for the purpose of simulations. The data generating process is $X(t)=\sum_{j=1}^{9}Z_{j}\phi_{j}(t)$, where $Z_{1},\ldots,Z_{9}]^{\top}$ is a 9-dimensional random vector with independent components. Suppose that $Z$ is a random variable following a Pareto distribution with tail index $\alpha=3$ and $N$ is a normal random variable with mean 0 and variance 0.5. We consider the following three cases for $Z_{1},\ldots,Z_{9}]^{\top}$:
1. $Z,N,N,N,\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{1}$.
2. $N,Z,N,N\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{2}$.
3. $N,N,Z,N\ldots,N]^{\top}$; the angular measure $\Gamma$ concentrates on $\phi_{3}$.
In all three cases, it follows from Proposition 7.1 and Example 7.3 of Meinguet & Segers (2010) that $X(t)$ is regularly varying with tail index $\alpha=3$. Fig. 1 displays simulated data with sample size of 100 for each of the three cases. The plots of simulated data clearly show that the angular measure $\Gamma$ represents the distribution of the shapes of extreme curves in that they are dominated by the shape of the functional axis $\phi_{j}$ on which $\Gamma$ concentrates.
@@ 3 Bivariate regular variation in Banach spaces
In order to describe the extremal dependence of two regularly varying random elements $X$ and $Y$ in $L^{2}\times L^{2}$, we need to identify their joint probabilistic behavior. We again study it in the more general space ${\mathbb{B}}^{2}$. We propose the following definition, which is a direct extension of Definition 2.1.
@@
A bivariate random element $X,Y]^{\top}$ in ${\mathbb{B}}^{2}$ is said to be jointly regularly varying with index $\alpha$, $\alpha>0$, if there exist a sequence $b(n)\to\infty$ and a measure $\mu$ in $M_{0}({\mathbb{B}}^{2})$ such that
$n{\rm pr}\left(\frac{(X,Y)}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\mu,\ \ \ \ n\to\infty.$ (4)
The joint exponent measure $\mu$ satisfies the scaling property
$\mu(tA)=t^{-\alpha}\mu(A),$ (5)
for any $t>0$ and Borel sets $A\subset{\mathbb{B}}_{0}^{2}:={\mathbb{B}}^{2}\setminus(\mbox{${0}$},\mbox{${0}$})$. Since $X$ and $Y$ are scaled by the same function $b(n)$, their marginal distributions are tail equivalent. As for a single Banach space, the measure $\mu$ is defined only up to a multiplicative constant, and the $b(n)$ are not unique. To introduce suitable normalizations, consider the sets
@@
According to Lemma A.1, the boundaries of these sets have $\mu$-measure zero. Thus, by normalizing $\mu$ using the usual condition
@@
we obtain that $n{\rm pr}(\|(X,Y)\|_{{\mathbb{B}}^{2}}>b(n))\to 1$.
Lemma 2.2 implies the following equivalent characterization of regular variation in ${\mathbb{B}}^{2}$.
Lemma 3.2.
A bivariate random element $X,Y]^{\top}$ in ${\mathbb{B}}^{2}$ is regularly varying with index $\alpha$, $\alpha>0$, if and only if there exists a sequence $b^{\prime}(n)\to\infty$ and a probability measure $\Gamma$ on ${\mathbb{S}}_{{\mathbb{B}}^{2}}$ (called the joint angular measure) such that for any $z>0$,
$n{\rm pr}\left(\|(X,Y)\|_{{\mathbb{B}}^{2}}>b^{\prime}(n)z,\ \frac{(X,Y)}{\|(X,Y)\|_{{\mathbb{B}}^{2}}}\in\cdot\right)\stackrel{{\scriptstyle w}}{{\longrightarrow}}cz^{-\alpha}\Gamma,\ \ \ \ n\to\infty,$ (8)
for some $c>0$.
With the scaling property (5) and normalization condition (7), we have $c=1$ in (8), and the joint angular measure $\Gamma$ on ${\mathbb{S}}_{{\mathbb{B}}^{2}}$ can be defined by
$\Gamma(S)=\mu\{(x,y)\in{\mathbb{B}}^{2}:\|(x,y)\|_{{\mathbb{B}}^{2}}>1,(x,y)/\|(x,y)\|_{{\mathbb{B}}^{2}}\in S\},$ (9)
for any Borel sets $S\subset{\mathbb{S}}_{{\mathbb{B}}^{2}}$. Throughout the paper, we use the norm on ${\mathbb{B}}^{2}$ defined by